OpenStudy (anonymous):

Identify which distributive property problem below would help you solve the following scenario: Your English class has 34 students. You would like to treat the class to cookies. You would like each person to have 3 cookies. Which expression would help you figure out how many cookies you would need to bake? 3(34 + 3) 3(20 - 4) 3(34 - 3) 3(30 + 4)

zepdrix (zepdrix):

Wow what a weird question, why in the world would you use the distributive property for something like this? So 34 students, 3 cookies each. 3(34) Of the options you are given, is there one that looks equivalent to 3(34), but written a little differently?

OpenStudy (anonymous):

ill try a

zepdrix (zepdrix):

a) 3(34+3) Hmm, 34+3 = 37 So that would simplify to 3(37). Hmmmmm, I don't think that's the one we want :)

OpenStudy (anonymous):

well ugh idk :/

zepdrix (zepdrix):

So they left the 3 alone. Try writing the number inside the brackets as 2 different numbers. Can you do that and make it equal one of the bracket choices? (34) = (20-4)? Hmm

OpenStudy (anonymous):

B !

zepdrix (zepdrix):

Hmm no :( (20-4)=(16)

zepdrix (zepdrix):

Lol your name is silly c: